Ingraham Inc, currently has $885,000 in accounts receivable, and its days sales outstanding (DSO) is 65 days


DSO AND ACCOUNTS RECEIVABLE 


Ingraham Inc, currently has $885,000 in accounts receivable, and its days sales outstanding (DSO) is 65 days. It wants to reduce its DSO to 35 days by pressuring more of its customers to pay their bills on time. If this policy is adopted, the company's average sales will fall by 15%. What will be the level of accounts receivable following the change? Assume a 365-day year. Do not round intermediate calculations. Round your answer to the nearest cent.


A firm has a profit margin of 6% and an equity multiplier of 2.9. Its sales are $50 million, and it has total assets of $25 million. What is its ROE? Do not round Intermediate calculations. Round your answer to two decimal places.


Hongland Corp's stock price at the end of last year was $48.50, and its book value per share was $25.00. What was its market/book ratio? 

 a. 1.55 

 b. 1.80 

 c. 1.63 

 d. 2.17 

 e. 1.94
Answer #1
Ingraham
Current DSO = ( Current account receivables / Current sales ) * Days in a year
65 = ( 885000 / Current sales ) * 365
Current sales = 885000 *365 / 65 4969615.38
Sales after change = Current sales * ( 1 - %fall ) = 4969615.38 * ( 1 - 15% ) 4224173.07
Level of accounts receivables following the change = DSO after change * Sales after change / Days in a year = 35 * 4224173.07 / 365    405057.69
DuPONT AND ROE
Equity multiplier = Total assets / Total equity
2.9 = 25000000 / Total equity
Total equity = 25000000 / 2.9 8620690
Net income = Sales * Profit margin = 50000000 * 6% 3000000
ROE = Net income / Total equity = 3000000 / 8620690 34.80%
Hoagland
Market/book ratio = Stock price per share / Book value per share = 48.50 / 25.00 1.94 Option e